Question 3 Nicole's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs $5.30 Nicole per pound, and type B coffee costs $4.05 per pound. This month, Nicole made 169 pounds of the blend, for a total cost of $803.20. How many pounds of type A coffee did she use?
Answer by addingup(3677)

Type A = 5.30/lb
Type B = 4.05/lb
--------------------------
1st equation
A + B = 169 lbs, subtract A from both sides:
B = 169-A
2nd equation
5.30A + 4.05B = 803.20
In this second equation, substitute for B with the result of the 1st equation:
5.30A + 4.05(169-A) = 803.2
5.30A + 684.45 - 4.05A = 803.2
1.25A = 118.75
A = 95
so there you have it, you need 95 pounds of coffee A.